OpenStudy (anonymous):

Suppose that in Sleepy Valley only 25% of those over 50 years old own CD players. Find the probabilty that among four randomly chosen people in that age group none owns a CD player.??

OpenStudy (sirm3d):

Prob of not owning a CD player = 1 - 0.25 = 0.75 all four people not owning a CD player = (0.75)^4

OpenStudy (anonymous):

I need your help with this question as well

OpenStudy (anonymous):

A stereo repair shop uses a two step diagnostic procedure to repair amplifiers. Step I locates the problem in the amplifier with probability 0.7 and Step II (which is executed only if step I fails to locate the problem) locates the problem with probability 0.7 . WHat is the probability that the procedure will fail to locate the problem?